OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (anonymous):

|2x + 6| − 4 = 20 |2x + 6|=24 |2x|=18 |x|=9 x=+9,-9

OpenStudy (anonymous):

for positive charge we can write 2x+6-4=20 =>2x+2=20 =>2x=20-2 =>x=18/2 =>x=9 for negative charge we can write -(2x+6)-4=20 =>-2x-6-4=20 =>-2x-10=20 =>-2x=20+10 =>-2x=30 =>x=-15